# On geometric spectral functionals

### v1

@@ 4 Chiral spectral functionals
An interesting generalization of the spectral functionals defined so far appears when one includes the grading $\chi$ for even spectral triples. We refer to them as chiral spectral functionals.
@@
$\mathscr{g}^{\chi}_{D}(u,w)=\mathscr{W}\!\mathrm{res}(\chi\hat{u}\hat{w}|D|^{-n}),\qquad u,w\in\Omega_{D}^{1},$ (4.1)
$\mathscr{G}_{D}^{\chi}(u,w)=\mathscr{W}\!\mathrm{res}\bigl{(}\chi\hat{u}\{D,\hat{w}\}D|D|^{-n}\bigr{)},\qquad u,w\in\Omega_{D}^{1},$ (4.2)
$\mathscr{T}_{D}^{\chi}(u,v,w)=\mathscr{W}\!\mathrm{res}(\chi\hat{u}\hat{v}\hat{w}DD^{-2m}),\qquad u,v,w\in\Omega^{1}_{D}.$ (4.3)
@@
$\mathscr{R}_{D}^{\chi}(f)=\mathscr{W}\!\mathrm{res}\bigl{(}\chi fD^{-2m+2}\bigr{)},\qquad f\in\mathcal{A}.$ (4.4)
In the following, we assume that the Dirac operator $D$ decomposes as $D=D_{0}+B$, where $D_{0}$ satisfies the assumptions outlined in the previous section and anticommutes with the grading $\chi$. Additionally, we require that the perturbation $B$ anticommutes with $\chi$ to ensure the anticommutativity of the full Dirac operator with this chiral structure. First, note that, similar to the non-chiral case, for $D=D_{0}+B$ satisfying the assumptions of this paper, the chiral metric functional satisfies $\mathscr{g}^{\chi}_{D}(u,w)=\mathscr{g}^{\chi}_{D_{0}}(u,w)$. Next, we have
@@
For $D\!=\!D_{0}+B$ satisfying the assumption of this paper, the chiral Einstein functional reads:
$\mathscr{G}_{D}^{\chi}(u,w)$ $=\mathscr{G}_{D_{0}}^{\chi}(u,w)$ (4.5)
$+\nu_{n-1}\int_{M}\mathrm{Tr}\,\chi\biggl{\{}iu_{a}w_{bc}\bigl{[}2(3-n)\gamma^{a}\delta^{bc}+2\delta^{ab}\gamma^{c}-2\delta^{ac}\gamma^{b}-(4-n)\gamma^{a}\gamma^{b}\gamma^{c}\bigr{]}B_{0}$
$\hskip 71.13188pt+iu_{a}w_{b}\bigl{[}(3-n)\gamma^{a}B_{b}-\gamma^{b}B_{a}+\delta^{ab}\gamma^{c}B_{c}-\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr{]}$
$\hskip 71.13188pt+\frac{1}{4}u_{a}w_{b}\bigl{(}[\gamma^{b},\gamma^{a}]B_{0}-2\gamma^{a}B_{0}\gamma^{b}\bigr{)}\gamma^{c}B_{0}\gamma^{c}\biggr{\}}.$
Proof 4.3.
The reasoning is analogous to the one in the proof of [3, Proposition II.1.], and we get
$\mathscr{G}_{D}^{\chi}(u,w)-\mathscr{G}_{D_{0}}^{\chi}(u,w)=$ (4.6)
$=\frac{\nu_{n-1}}{24}\int\hbox{Tr\ }\Big{(}24\gamma u_{a}\gamma^{a}(i\gamma^{c}\gamma^{b}w_{bc}B_{0}\!+i\!\gamma^{c}\gamma^{b}w_{b}B_{c}+w_{b}B_{0}\gamma^{b}B_{0})$
$\;\;+12i\gamma\bigl{[}-u_{a}w_{bd}\gamma^{a}\gamma^{d}\gamma^{b}\gamma^{c}\!+\!iu_{a}w_{b}\gamma^{a}(2\delta^{bc}B_{0}\!+\!\{\gamma^{b},B_{0}\}\gamma^{c}\!-\!\gamma^{b}\{\gamma^{c},B_{0}\})\bigr{]}\{\gamma^{c},B_{0}\}$
$\;\;+6\gamma(u_{c}w_{b}\gamma^{c}\gamma^{a}\!+\!u_{c}w_{a}\gamma^{c}\gamma^{b}\!-\!u_{c}w_{d}\gamma^{c}\gamma^{d}\delta_{ab})\bigl{[}-2i\{\gamma^{a},B_{b}\}\!+\!\{\gamma^{a},B_{0}\}\{\gamma^{b},B_{0}\}\bigr{]}\Big{)}.$
First we collect the terms with $w_{bc}$
$\frac{i}{2}u_{a}w_{bc}\mathrm{Tr}\,\gamma[2\gamma^{a}\gamma^{c}\gamma^{b}B_{0}-\gamma^{a}\gamma^{c}\gamma^{b}\gamma^{d}\{\gamma^{d},B_{0}\}]$ (4.7)
$=\frac{i}{2}u_{a}w_{bc}\hbox{Tr\ }\gamma[(2-n)\gamma^{a}\gamma^{c}\gamma^{b}B_{0}-\gamma^{a}\gamma^{c}\gamma^{b}\gamma^{d}B_{0}\gamma^{d}]$
$=\frac{i}{2}u_{a}w_{bc}\mathrm{Tr}\,[(2-n)\gamma\gamma^{a}\gamma^{c}\gamma^{b}B_{0}+2\delta^{ad}\gamma\gamma^{c}\gamma^{b}\gamma^{d}B_{0}-2\delta^{cd}\gamma\gamma^{a}\gamma^{b}\gamma^{d}B_{0}$
$\hskip 50.0pt+2\delta^{bd}\gamma\gamma^{a}\gamma^{c}\gamma^{d}B_{0}-n\gamma\gamma^{a}\gamma^{c}\gamma^{b}B_{0}]$
$=iu_{a}w_{bc}\mathrm{Tr}\,[(2-n)\gamma\gamma^{a}\gamma^{c}\gamma^{b}B_{0}+\gamma\gamma^{c}\gamma^{b}\gamma^{a}B_{0}-\gamma\gamma^{a}\gamma^{b}\gamma^{c}B_{0}]$
$=iu_{a}w_{bc}\mathrm{Tr}\bigl{[}(4-n)\gamma\gamma^{a}\gamma^{c}\gamma^{b}B_{0}+2\gamma(\delta^{ab}\gamma^{c}-\delta^{ac}\gamma^{b}-\delta^{bc}\gamma^{a})B_{0}\bigr{]}$
$=iu_{a}w_{bc}\mathrm{Tr}\bigl{[}\gamma\bigl{(}(6-2n)\gamma^{a}\delta^{bc}+2\delta^{ab}\gamma^{c}-2\delta^{ac}\gamma^{b}-(4-n)\gamma^{a}\gamma^{b}\gamma^{c}\bigr{)}B_{0}\bigr{]}$
@@
$\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}\gamma\gamma^{a}(2\gamma^{c}\gamma^{b}B_{c}-\gamma^{c}\{\gamma^{c},B_{b}\}-\gamma^{c}\{\gamma^{b},B_{c}\}+\gamma^{b}\{\gamma^{c},B_{c}\})\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}\gamma\gamma^{a}(\gamma^{c}\gamma^{b}B_{c}-nB_{b}-\gamma^{c}B_{b}\gamma^{c}-\gamma^{c}B_{c}\gamma^{b}+\gamma^{b}\gamma^{c}B_{c}+\gamma^{b}B_{c}\gamma^{c})\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}\gamma\gamma^{a}\{\gamma^{c},\gamma^{b}\}B_{c}-n\gamma\gamma^{a}B_{b}-\gamma\gamma^{a}\gamma^{c}B_{b}\gamma^{c}-\gamma\gamma^{a}\gamma^{c}B_{c}\gamma^{b}+\gamma\gamma^{a}\gamma^{b}B_{c}\gamma^{c}\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}(2-n)\gamma\gamma^{a}B_{b}+\gamma(2\delta^{ac}-\gamma^{a}\gamma^{c})\gamma^{c}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}+\gamma\gamma^{a}\gamma^{b}B_{b}\gamma^{c}\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}(2-n)\gamma\gamma^{a}B_{b}+2\gamma\gamma^{a}B_{b}-n\gamma\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-\gamma\gamma^{c}\gamma^{a}\gamma^{b}B_{c}\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}(2-n)\gamma\gamma^{a}B_{b}+2\gamma\gamma^{a}B_{b}-n\gamma\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-\gamma(2\delta^{ac}-\gamma^{a}\gamma^{c})\gamma^{b}B_{c}\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}2(2-n)\gamma\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-2\gamma\gamma^{b}B_{a}+\gamma\gamma^{a}(2\delta^{cb}-\gamma^{b}\gamma^{c})B_{c}\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}2(2-n)\gamma\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-2\gamma\gamma^{b}B_{a}+2\gamma\gamma^{a}B_{b}-\gamma\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr{]}$
$=\frac{i}{2}u_{a}w_{b}\mathrm{Tr}\,\bigl{[}2(3-n)\gamma\gamma^{a}B_{b}+\gamma(2\delta^{ab}-\gamma^{a}\gamma^{b})\gamma^{c}B_{c}-2\gamma\gamma^{b}B_{a}-\gamma\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr{]}$
$=iu_{a}w_{b}\mathrm{Tr}\,\bigl{[}(3-n)\gamma\gamma^{a}B_{b}-\gamma\gamma^{b}B_{a}+\delta^{ab}\gamma\gamma^{c}B_{c}-\gamma\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr{]}.$ (4.8)
@@
$\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\,\chi\gamma^{a}\Bigl{[}4B_{0}\gamma^{b}B_{0}-4B_{0}\{\gamma^{b},B_{0}\}-2\{\gamma^{b},B_{0}\}\gamma^{c}\{\gamma^{c},B_{0}\}+2\gamma^{b}\{\gamma^{c},B_{0}\}^{2}$
$\hskip 50.0pt+\gamma^{c}\{\gamma^{c},B_{0}\}\{\gamma^{b},B_{0}\}+\gamma^{c}\{\gamma^{b},B_{0}\}\{\gamma^{c},B_{0}\}-\gamma^{b}\{\gamma^{c},B_{0}\}^{2}\Bigr{]}$
$=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\,\chi\gamma^{a}\Bigl{[}4B_{0}\gamma^{b}B_{0}-4B_{0}\gamma^{b}B_{0}-4B_{0}^{2}\gamma^{b}+\gamma^{c}\{\gamma^{c},B_{0}\}\gamma^{b}B_{0}+\gamma^{c}\{\gamma^{c},B_{0}\}B_{0}\gamma^{b}$
$\hskip 78.24507pt+\bigl{(}-2\gamma^{b}B_{0}\gamma^{c}-2B_{0}\gamma^{b}\gamma^{c}+2\gamma^{b}\gamma^{c}B_{0}+2\gamma^{b}B_{0}\gamma^{c}+\gamma^{c}\gamma^{b}B_{0}+\gamma^{c}B_{0}\gamma^{b}$
$\hskip 256.0748pt-\gamma^{b}\gamma^{c}B_{0}-\gamma^{b}B_{0}\gamma^{c}\bigl{)}\{\gamma^{c},B_{0}\}\Bigr{]}$
$=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\,\chi\Bigl{(}4\gamma^{b}\gamma^{a}B_{0}^{2}+n\gamma^{a}B_{0}\gamma^{b}B_{0}+\gamma^{a}\gamma^{c}B_{0}\gamma^{c}\gamma^{b}B_{0}-n\gamma^{b}\gamma^{a}B_{0}^{2}-\gamma^{b}\gamma^{a}\gamma^{c}B_{0}\gamma^{c}B_{0}$
$\hskip 71.13188pt+2\gamma^{a}B_{0}\{\gamma^{b},B_{0}\}-2n\gamma^{a}B_{0}\gamma^{b}B_{0}-2\gamma^{a}B_{0}\gamma^{b}\gamma^{c}B_{0}\gamma^{c}+\gamma^{a}\gamma^{c}B_{0}\gamma^{b}\gamma^{c}B_{0}$
$\hskip 71.13188pt+\gamma^{a}\gamma^{c}B_{0}\gamma^{b}B_{0}\gamma^{c}-n\gamma^{a}\gamma^{b}B_{0}^{2}-\gamma^{a}\gamma^{b}B_{0}\gamma^{c}B_{0}\gamma^{c}\Bigr{)}$
$=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\,\chi\bigl{[}(2-n)\gamma^{a}\gamma^{b}B_{0}^{2}+(2-n)\gamma^{b}\gamma^{a}B_{0}^{2}+(2-n)\gamma^{a}B_{0}\gamma^{b}B_{0}-\gamma^{b}\gamma^{a}\gamma^{c}B_{0}\gamma^{c}B_{0}$
$\hskip 71.13188pt+4B_{0}\gamma^{b}\gamma^{a}B_{0}-2\gamma^{a}\gamma^{c}B_{0}\gamma^{b}\gamma^{c}B_{0}+(n-2)\gamma^{a}B_{0}\gamma^{b}B_{0}-\gamma^{a}\gamma^{b}B_{0}\gamma^{c}B_{0}\gamma^{c}\bigr{]}$
$=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\,\chi\bigl{[}2(2-n)\delta^{ab}B_{0}^{2}-2\gamma^{b}B_{0}\gamma^{a}B_{0}+2\gamma^{a}B_{0}\gamma^{b}B_{0}+\gamma^{b}\gamma^{a}B_{0}\gamma^{c}B_{0}\gamma^{c}+4B_{0}\gamma^{b}\gamma^{a}B_{0}$
$\hskip 71.13188pt-4B_{0}\gamma^{b}\gamma^{a}B_{0}-2\gamma^{a}B_{0}\gamma^{b}\gamma^{c}B_{0}\gamma^{c}-\gamma^{a}\gamma^{b}B_{0}\gamma^{c}B_{0}\gamma^{c}\bigr{]}$
$=\frac{1}{4}u_{q}w_{b}\mathrm{Tr}\,\chi\bigl{(}[\gamma^{b},\gamma^{a}]B_{0}-2\gamma^{a}B_{0}\gamma^{b}\bigr{)}\gamma^{c}B_{0}\gamma^{c}.$ (4.9)
@@
For $D\!=\!D_{0}+B$ satisfying the assumption of this paper, the chiral torsion functional reads:
$\mathscr{T}^{\chi}_{D}(u,v,w)=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\left[\chi\hat{u}\hat{v}\hat{w}(2B_{0}-\gamma^{a}\{\gamma^{a},B_{0}\})\right].$ (4.10)
Proof 4.5.
It follows from Proposition 2.5.
@@
For $D\!=\!D_{0}+B$ satisfying the assumption of this paper, the chiral scalar curvature functional reads:
$\mathscr{R}^{\chi}_{D}(f)=\mathscr{R}^{\chi}_{D_{0}}(f)-i\frac{\nu_{n-1}}{2}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\,\bigl{(}\chi\gamma^{a}B_{a}\bigr{)}.$ (4.11)
@@
$\mathscr{W}\!\mathrm{res}(\chi D^{-n+2})=\frac{\nu_{n-1}}{24}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\,\bigl{[}\chi(-12Q-2R+6P_{aa}-3S_{a}S_{a})\bigr{]},$ (4.12)

### v2

@@ 4 Chiral spectral functionals
An interesting generalization of the spectral functionals defined so far appears when one includes the grading $\chi$ for even spectral triples. We refer to them as chiral spectral functionals. Their geometrical meaning is rather simple, as they provide pseudoscalar quantities instead of scalar ones. This extends the realm of possible functionals and, as we shall discuss in this section for the spin manifolds, allows us to represent some obvious pseudoscalar functionals as chiral spectral functionals. It remains to be studied in more detail whether some functionals of this type can capture, for example, other perturbations of the Dirac operator that arise from torsion.
@@
$\mathscr{g}^{\chi}_{D}(u,w)=\mathscr{W}\mathrm{res}(\chi\hat{u}\hat{w}|D|^{-n}),\qquad u,w\in\Omega_{D}^{1},$
[…]
$\mathscr{T}_{D}^{\chi}(u,v,w)=\mathscr{W}\mathrm{res}\bigl(\chi\hat{u}\hat{v}\hat{w}DD^{-2m}\bigr),\qquad u,v,w\in\Omega^{1}_{D},$
@@
$\mathscr{R}_{D}^{\chi}(f)=\mathscr{W}\mathrm{res}\bigl(\chi fD^{-2m+2}\bigr),\qquad f\in\mathcal{A}.$
In the following, we assume that the Dirac operator $D$ decomposes as $D=D_{0}+B$, where $D_{0}$ satisfies the assumptions outlined in the previous section and anticommutes with the grading $\chi$. Additionally, we require that the perturbation $B$ anticommutes with $\chi$ to ensure the anticommutativity of the full Dirac operator with this chiral structure. First, note that, similar to the non-chiral case, for $D=D_{0}+B$ satisfying the assumptions of this paper, the chiral metric functional satisfies $\mathscr{g}^{\chi}_{D}(u,w)=\mathscr{g}^{\chi}_{D_{0}}(u,w)$. Next, we have the following.
@@
For $D=D_{0}+B$ satisfying the assumption of this paper, the chiral Einstein functional reads
$\mathscr{G}_{D}^{\chi}(u,w)={}$ $\mathscr{G}_{D_{0}}^{\chi}(u,w)+\nu_{n-1}\int_{M}\mathrm{Tr}\chi\Bigr\{{\rm i}u_{a}w_{bc}\bigl[2(3-n)\gamma^{a}\delta^{bc}+2\delta^{ab}\gamma^{c}-2\delta^{ac}\gamma^{b}$
$-(4-n)\gamma^{a}\gamma^{b}\gamma^{c}\bigr]B_{0}+{\rm i}u_{a}w_{b}\bigl[(3-n)\gamma^{a}B_{b}-\gamma^{b}B_{a}+\delta^{ab}\gamma^{c}B_{c}-\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr]$
$+\frac{1}{4}u_{a}w_{b}\bigl(\bigl[\gamma^{b},\gamma^{a}\bigr]B_{0}-2\gamma^{a}B_{0}\gamma^{b}\bigr)\gamma^{c}B_{0}\gamma^{c}\Bigr\}.$
Proof.
The reasoning is analogous to the one in the proof of [2, Proposition II.1.], and we get
$\mathscr{G}_{D}^{\chi}(u,w)-\mathscr{G}_{D_{0}}^{\chi}(u,w)$
$\qquad=\frac{\nu_{n-1}}{24}\int\operatorname{Tr}\bigl(24\chi u_{a}\gamma^{a}\bigl({\rm i}\gamma^{c}\gamma^{b}w_{bc}B_{0}+{\rm i}\gamma^{c}\gamma^{b}w_{b}B_{c}+w_{b}B_{0}\gamma^{b}B_{0}\bigr)$
$\phantom{\qquad=}{}+12{\rm i}\chi\bigl[-u_{a}w_{bd}\gamma^{a}\gamma^{d}\gamma^{b}\gamma^{c}+{\rm i}u_{a}w_{b}\gamma^{a}(2\delta^{bc}B_{0}+\bigl\{\gamma^{b},B_{0}\bigr\}\gamma^{c}-\gamma^{b}\{\gamma^{c},B_{0}\})\bigr]\{\gamma^{c},B_{0}\}$
$\phantom{\qquad=}{}+6\chi\bigl(u_{c}w_{b}\gamma^{c}\gamma^{a}+u_{c}w_{a}\gamma^{c}\gamma^{b}-u_{c}w_{d}\gamma^{c}\gamma^{d}\delta_{ab}\bigr)\bigl[-2{\rm i}\{\gamma^{a},B_{b}\}+\bigl\{\gamma^{a},B_{0}\bigr\}\bigl\{\gamma^{b},B_{0}\bigr\}\bigr]\bigr).$
First, we collect the terms with $w_{bc}$
$\frac{\rm i}{2}u_{a}w_{bc}\mathrm{Tr}\chi\bigl[2\gamma^{a}\gamma^{c}\gamma^{b}B_{0}-\gamma^{a}\gamma^{c}\gamma^{b}\gamma^{d}\bigl\{\gamma^{d},B_{0}\bigr\}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{bc}\operatorname{Tr}\chi\bigl[(2-n)\gamma^{a}\gamma^{c}\gamma^{b}B_{0}-\gamma^{a}\gamma^{c}\gamma^{b}\gamma^{d}B_{0}\gamma^{d}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{bc}\mathrm{Tr}\bigl[(2-n)\chi\gamma^{a}\gamma^{c}\gamma^{b}B_{0}+2\delta^{ad}\chi\gamma^{c}\gamma^{b}\gamma^{d}B_{0}-2\delta^{cd}\chi\gamma^{a}\gamma^{b}\gamma^{d}B_{0}$
$\phantom{\qquad=}{}+2\delta^{bd}\chi\gamma^{a}\gamma^{c}\gamma^{d}B_{0}-n\chi\gamma^{a}\gamma^{c}\gamma^{b}B_{0}\bigr]$
$\qquad={\rm i}u_{a}w_{bc}\mathrm{Tr}\bigl[(2-n)\chi\gamma^{a}\gamma^{c}\gamma^{b}B_{0}+\chi\gamma^{c}\gamma^{b}\gamma^{a}B_{0}-\chi\gamma^{a}\gamma^{b}\gamma^{c}B_{0}\bigr]$
$\qquad={\rm i}u_{a}w_{bc}\mathrm{Tr}\bigl[(4-n)\chi\gamma^{a}\gamma^{c}\gamma^{b}B_{0}+2\chi\bigl(\delta^{ab}\gamma^{c}-\delta^{ac}\gamma^{b}-\delta^{bc}\gamma^{a}\bigr)B_{0}\bigr]$
$\qquad={\rm i}u_{a}w_{bc}\mathrm{Tr}\bigl[\chi\bigl((6-2n)\gamma^{a}\delta^{bc}+2\delta^{ab}\gamma^{c}-2\delta^{ac}\gamma^{b}-(4-n)\gamma^{a}\gamma^{b}\gamma^{c}\bigr)B_{0}\bigr].$
@@
$\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[\chi\gamma^{a}\bigl(2\gamma^{c}\gamma^{b}B_{c}-\gamma^{c}\{\gamma^{c},B_{b}\}-\gamma^{c}\{\gamma^{b},B_{c}\}+\gamma^{b}\{\gamma^{c},B_{c}\}\bigr)\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[\chi\gamma^{a}\bigl(\gamma^{c}\gamma^{b}B_{c}-nB_{b}-\gamma^{c}B_{b}\gamma^{c}-\gamma^{c}B_{c}\gamma^{b}+\gamma^{b}\gamma^{c}B_{c}+\gamma^{b}B_{c}\gamma^{c}\bigr)\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[\chi\gamma^{a}\bigl\{\gamma^{c},\gamma^{b}\bigr\}B_{c}-n\gamma\gamma^{a}B_{b}-\gamma\gamma^{a}\gamma^{c}B_{b}\gamma^{c}-\gamma\gamma^{a}\gamma^{c}B_{c}\gamma^{b}+\gamma\gamma^{a}\gamma^{b}B_{c}\gamma^{c}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[(2-n)\chi\gamma^{a}B_{b}+\gamma(2\delta^{ac}-\gamma^{a}\gamma^{c})\gamma^{c}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}+\gamma\gamma^{a}\gamma^{b}B_{b}\gamma^{c}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[(2-n)\chi\gamma^{a}B_{b}+2\gamma\gamma^{a}B_{b}-n\gamma\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-\gamma\gamma^{c}\gamma^{a}\gamma^{b}B_{c}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[(2-n)\chi\gamma^{a}B_{b}+2\gamma\gamma^{a}B_{b}-n\gamma\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-\gamma(2\delta^{ac}-\gamma^{a}\gamma^{c})\gamma^{b}B_{c}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[2(2-n)\chi\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-2\gamma\gamma^{b}B_{a}+\gamma\gamma^{a}\bigl(2\delta^{cb}-\gamma^{b}\gamma^{c}\bigr)B_{c}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[2(2-n)\chi\gamma^{a}B_{b}+\gamma\gamma^{b}\gamma^{a}\gamma^{c}B_{c}-2\gamma\gamma^{b}B_{a}+2\gamma\gamma^{a}B_{b}-\gamma\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr]$
$\qquad=\frac{\rm i}{2}u_{a}w_{b}\mathrm{Tr}\bigl[2(3-n)\chi\gamma^{a}B_{b}+\gamma\bigl(2\delta^{ab}-\gamma^{a}\gamma^{b}\bigr)\gamma^{c}B_{c}-2\gamma\gamma^{b}B_{a}-\gamma\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr]$
$\qquad={\rm i}u_{a}w_{b}\mathrm{Tr}\bigl[(3-n)\chi\gamma^{a}B_{b}-\gamma\gamma^{b}B_{a}+\delta^{ab}\gamma\gamma^{c}B_{c}-\gamma\gamma^{a}\gamma^{b}\gamma^{c}B_{c}\bigr].$
@@
$\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\chi\gamma^{a}\bigl[4B_{0}\gamma^{b}B_{0}-4B_{0}\bigl\{\gamma^{b},B_{0}\bigr\}-2\bigl\{\gamma^{b},B_{0}\bigr\}\gamma^{c}\{\gamma^{c},B_{0}\}+2\gamma^{b}\{\gamma^{c},B_{0}\}^{2}$
$\qquad+\gamma^{c}\{\gamma^{c},B_{0}\}\bigl\{\gamma^{b},B_{0}\bigr\}+\gamma^{c}\bigl\{\gamma^{b},B_{0}\bigr\}\{\gamma^{c},B_{0}\}-\gamma^{b}\{\gamma^{c},B_{0}\}^{2}\bigr]$
$\phantom{\qquad+}{}=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\chi\gamma^{a}\bigl[4B_{0}\gamma^{b}B_{0}-4B_{0}\gamma^{b}B_{0}-4B_{0}^{2}\gamma^{b}+\gamma^{c}\{\gamma^{c},B_{0}\}\gamma^{b}B_{0}$
$\phantom{\qquad+=}{}+\gamma^{c}\{\gamma^{c},B_{0}\}B_{0}\gamma^{b}+\bigl(-2\gamma^{b}B_{0}\gamma^{c}-2B_{0}\gamma^{b}\gamma^{c}+2\gamma^{b}\gamma^{c}B_{0}+2\gamma^{b}B_{0}\gamma^{c}$
$\phantom{\qquad+=}{}+\gamma^{c}\gamma^{b}B_{0}+\gamma^{c}B_{0}\gamma^{b}-\gamma^{b}\gamma^{c}B_{0}-\gamma^{b}B_{0}\gamma^{c}\bigr)\{\gamma^{c},B_{0}\}\bigr]$
$\phantom{\qquad+}{}=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\chi\bigl(4\gamma^{b}\gamma^{a}B_{0}^{2}+n\gamma^{a}B_{0}\gamma^{b}B_{0}+\gamma^{a}\gamma^{c}B_{0}\gamma^{c}\gamma^{b}B_{0}-n\gamma^{b}\gamma^{a}B_{0}^{2}$
$\phantom{\qquad+=}{}-\gamma^{b}\gamma^{a}\gamma^{c}B_{0}\gamma^{c}B_{0}+2\gamma^{a}B_{0}\bigl\{\gamma^{b},B_{0}\bigr\}-2n\gamma^{a}B_{0}\gamma^{b}B_{0}-2\gamma^{a}B_{0}\gamma^{b}\gamma^{c}B_{0}\gamma^{c}$
$\phantom{\qquad+=}{}+\gamma^{a}\gamma^{c}B_{0}\gamma^{b}\gamma^{c}B_{0}+\gamma^{a}\gamma^{c}B_{0}\gamma^{b}B_{0}\gamma^{c}-n\gamma^{a}\gamma^{b}B_{0}^{2}-\gamma^{a}\gamma^{b}B_{0}\gamma^{c}B_{0}\gamma^{c}\bigr)$
$\phantom{\qquad+}{}=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\chi\bigl[(2-n)\gamma^{a}\gamma^{b}B_{0}^{2}+(2-n)\gamma^{b}\gamma^{a}B_{0}^{2}+(2-n)\gamma^{a}B_{0}\gamma^{b}B_{0}$
$\phantom{\qquad+=}{}-\gamma^{b}\gamma^{a}\gamma^{c}B_{0}\gamma^{c}B_{0}+4B_{0}\gamma^{b}\gamma^{a}B_{0}-2\gamma^{a}\gamma^{c}B_{0}\gamma^{b}\gamma^{c}B_{0}$
$\phantom{\qquad+=}{}+(n-2)\gamma^{a}B_{0}\gamma^{b}B_{0}-\gamma^{a}\gamma^{b}B_{0}\gamma^{c}B_{0}\gamma^{c}\bigr]$
$\phantom{\qquad+}{}=\frac{1}{4}u_{a}w_{b}\mathrm{Tr}\chi\bigl[2(2-n)\delta^{ab}B_{0}^{2}-2\gamma^{b}B_{0}\gamma^{a}B_{0}+2\gamma^{a}B_{0}\gamma^{b}B_{0}+\gamma^{b}\gamma^{a}B_{0}\gamma^{c}B_{0}\gamma^{c}$
$\phantom{\qquad+=}{}+4B_{0}\gamma^{b}\gamma^{a}B_{0}-4B_{0}\gamma^{b}\gamma^{a}B_{0}-2\gamma^{a}B_{0}\gamma^{b}\gamma^{c}B_{0}\gamma^{c}-\gamma^{a}\gamma^{b}B_{0}\gamma^{c}B_{0}\gamma^{c}\bigr]$
$\phantom{\qquad+}{}=\frac{1}{4}u_{q}w_{b}\mathrm{Tr}\chi\bigl(\bigl[\gamma^{b},\gamma^{a}\bigr]B_{0}-2\gamma^{a}B_{0}\gamma^{b}\bigr)\gamma^{c}B_{0}\gamma^{c}.$
@@
For $D=D_{0}+B$ satisfying the assumption of this paper, the chiral torsion functional reads
$\mathscr{T}^{\chi}_{D}(u,v,w)=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\left[\chi\hat{u}\hat{v}\hat{w}(2B_{0}-\gamma^{a}\{\gamma^{a},B_{0}\})\right].$
Proof.
It follows from Proposition 2.5. ∎
@@
For $D=D_{0}+B$ satisfying the assumption of this paper, the chiral scalar curvature functional reads
$\mathscr{R}^{\chi}_{D}(f)=\mathscr{R}^{\chi}_{D_{0}}(f)-{\rm i}\frac{\nu_{n-1}}{2}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}(\chi\gamma^{a}B_{a}).$
@@
$\mathscr{W}\mathrm{res}\bigl(\chi D^{-n+2}\bigr)=\frac{\nu_{n-1}}{24}(n-2)\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl[\chi(-12Q-2R+6P_{aa}-3S_{a}S_{a})\bigr],$